OpenStudy (anonymous):

find the lcd 5\x-6' 7\x^2-12x+36

OpenStudy (anonymous):

\[\frac{5}{x-6}, \frac{7}{x^2-12x+36}\]\[\frac{5}{x-6}, \frac{7}{(x-6)(x-6)}\]\[\frac{5}{x-6} \bullet \frac{(x-6)}{(x-6)}, \frac{7}{(x-6)(x-6)}\]\[\frac{5x-30}{(x-6)(x-6)}, \frac{7}{(x-6)(x-6)}\]

OpenStudy (anonymous):

ok